OpenStudy (anonymous):

1. Jeremiah is making a shake by mixing two different protein powders, measured in ounces. The strawberry-flavored powder has 4 grams of protein per ounce. The banana-flavored powder has 3 grams of protein per ounce. He wants the drink to have a total of 6 ounces of powder and contain 22 grams of protein. a.Create a system of linear equations to represent the situation that will determine the exact number of ounces needed for each power. b.Based on the system you wrote, which algebraic method of solving systems of equations will you choose to solve the system? Explain your choice. c.Solv

OpenStudy (anonymous):

c.Solve the system algebraically, showing your work. Then interpret the solution by explaining what your solution represents in the context of the problem.(sorry bout the cut off)

OpenStudy (anonymous):

S + B = 6 ounces of powder 4 S + 3 B = 22 grams of protein solve multiply first equation by -4 for each term, both sides of equation add the equations, getting rid of S that way - B = -24 + 22 = - 2 B = 2 then solve for S

OpenStudy (anonymous):

@douglaswinslowcooper Thank you so much:))

OpenStudy (anonymous):

Glad to have helped!

OpenStudy (anonymous):

Thanks hehe:))